OpenStudy (anonymous):

Solve The Equation Using The Zero-Product Property. (x+9)(x+2)=0

OpenStudy (anonymous):

x+9=0 and x+2=0 solve it from there

OpenStudy (anonymous):

would it be x+7=0? @jstew2344

OpenStudy (anonymous):

this equation will have 2 answers for x. The first answer is x+9=0 which is x=-9 and the second answer is x+2=0 which is x=-2 so your answers are x=-2 and x=-9

OpenStudy (anonymous):

lol wow i didnt even think of that to be honest but thanks so much @jstew2344